OpenStudy (anonymous):

Add negative two and three-fifths plus negative three and one-half. negative nine-tenths negative six and one-tenth negative four and one-tenth negative one and one-tenth

OpenStudy (tkhunny):

What's preventing you from doing this? \(-2\dfrac{3}{5} + \left(-3\dfrac{1}{2}\right) = -2\dfrac{3}{5} - 3\dfrac{1}{2} = -\dfrac{13}{5} - \dfrac{7}{2}\) Make sure you get the signs right. Generally, get rid of the mixed fractions. Find a common denominator. Go!!